OpenStudy (anonymous):

Simplify four to the ninth power times four to the negative seventh power . four to the negative sixteenth power negative four to the second power four to the negative sixty-third power four to the second power

OpenStudy (anonymous):

this is what it looks like in math... \[4^{9} \times 4^{-7}\]

OpenStudy (anonymous):

4^2 is your answer

OpenStudy (anonymous):

He needs it simplified, not solved.

OpenStudy (anonymous):

oh wait, my bad. night-watcher is right.

OpenStudy (paki):

4^9*4^-7 = 4^2 = 4*4 = 16

OpenStudy (anonymous):

Hahah thank you guys c:

OpenStudy (anonymous):

No prob.

OpenStudy (paki):

pleasure

OpenStudy (anonymous):

No problem